OpenStudy (anonymous):

Two problems I need help on: Given the point (2, 3) and the slope of 4, find y when x = 22. 78 83 88 91 And Given the function f(x) = 4(x+3) − 5, solve for the inverse function when x = 3. −21 −16 −10 −1

hero (hero):

y = mx + b 3 = 4(2) + b 3 = 8 + b 3 - 8 = b -5 = b So m = 4 and b = -5 so y = 4x - 5 When x = 22: y = 4(22) - 5 y = 88 - 5 y = 83

OpenStudy (anonymous):

Oh okay... so would that work for the second one, even though its asking for the inverse?

OpenStudy (anonymous):

oh wait no...because its not the same set up, it already has the equations set up

OpenStudy (anonymous):

so would I do (3)=4(3-3)-5 Because its asking for the inverse?

OpenStudy (anonymous):

or (3)=4(3-3)+5

hero (hero):

Okay, yeah, you have to find the inverse first.

hero (hero):

f(x) = 4(x + 3) - 5 Change f(x) to y: y = 4(x + 3) - 5 Swap x and y: x = 4(y + 3) - 5 Isolate y: x + 5 = 4(y + 3) \(\dfrac{x + 5}{4} = y + 3\) Now input x = 3: \(\dfrac{x + 5}{4} - 3 = y\) \(\dfrac{3 + 5}{4} - 3= y \)

OpenStudy (anonymous):

okay I'm doing something wrong...I keep getting 1.25, but that's not a option

hero (hero):

\(\dfrac{8}{4} - 3 = y\) \(2 - 3 = y\) \(-1 = y\)

OpenStudy (anonymous):

Okay, just got that...my calculator was being dumb! Thanks ^-^

hero (hero):

You should be able to do it without a calculator.